Suppose that demand and supply are given by the following equations: P 85-18.Q P = 30+4 Q What is the equilibrium price?
Answer #1

Answer
the equilibrium is at Qd=Qs
85-18Q=30+4Q
22Q=55
Q=2.5
P=85-18*2.5=40

the price is $40